OpenStudy (ilovebmth1234):

What is the energy of a photon of light having a wavelength of 4.5 x 10-7 meters? (Planck’s constant is 6.626 x 10-34 joule seconds; the speed of light is 2.998 x 108 m/s) 2.9817 x 10-40 joules 4.4143 x 10-19 joules 6.6745 x 10-26 joules second 9.9456 x 10-49 joules 19.864 x 10-26 joules/second

OpenStudy (ilovebmth1234):

@Missiey im sorry i know you helped me a lot yesterday but you think you could help me one or two more times xD

OpenStudy (anonymous):

yes give me a sec

OpenStudy (ilovebmth1234):

thank you so much!

OpenStudy (anonymous):

6.6745 x 10-26 joules secon is what i got

OpenStudy (anonymous):

u have another one

OpenStudy (ilovebmth1234):

yes, The frequency of a given region of the electromagnetic spectrum is more than 3 x 1019 HZ. Note that the speed of light is 2.998 x 108 m/s. Which waves are found in this region? gamma rays visible light microwaves radio waves x–rays

OpenStudy (ilovebmth1234):

my guess was gamma rays

OpenStudy (anonymous):

The wavelength is L=c/f = 2.998*10^8/3*10^19 = ~10^-11m = 0.01nm That is border-line between gamma and x-ray, but if you'd have to choose just one answer based on 10^-11m wavelength as threshold it would be gamma as the result of the calculation is slightly below 10^-11 2.998*10^8/3*10^19 = 0.9993*10^-11 in short yes gamma ray

OpenStudy (anonymous):

sorry wanted to atleast explain alittle

OpenStudy (ilovebmth1234):

thats okay lol explaining is good :) thanks

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

anymore

OpenStudy (ilovebmth1234):

maybe one i have to answer this question i know this one(:

OpenStudy (ilovebmth1234):

In what form does all radiant energy travel through space? In the form of light. In the form of electromagnetic radiation. In the form of accelerating protons. In the form of heat energy. In the form of accelerating neutrons

OpenStudy (anonymous):

Radiant belongs to the electromagnetic force and travels through the electromagnetic field.

OpenStudy (ilovebmth1234):

so B

OpenStudy (anonymous):

yes

OpenStudy (ilovebmth1234):

thank you so much for all your help!

OpenStudy (anonymous):

anytime